OpenStudy (anonymous):

The school board announced they are increasing all athletic fees by a fixed percent. A formula for determining the new fee is as follows: c= b-a ----- <- (this is a fraction btw, just very poorly shown) a where b= new athletic fee a= old athletic fee c= percent increase If the new athletic fee is $22, and the old fee is $20. How much is the percent increase in fees? Write your answer in a percent, please.

OpenStudy (mathmate):

Use the given formula: a=20, b=22 \(c=\frac{b-a}{a}=\frac{22-20}{20}\times 100\%=?\)

OpenStudy (xapproachesinfinity):

mathmate already give you the answer? what is the problem

OpenStudy (xapproachesinfinity):

what is 22-20?